OpenStudy (anonymous):

Find derivative of the function: g(x) = ln(e^x)/(1 + e^x)

myininaya (myininaya):

is that just natural log of the top or the whole thing?

OpenStudy (anonymous):

natural log of the whole thing

OpenStudy (anonymous):

simplify first:\[\ln (e^{x}) = x\] now we have:\[x \over (1+e^{x})\] use quotien rule to find the derivative

myininaya (myininaya):

g(x)=ln(e^x)-ln(1+e^x) g(x)=xlne-ln(1+e^x) g'(x)=lne-(e^x)/(1+e^x)

OpenStudy (anonymous):

oh nvm my answer then..

OpenStudy (anonymous):

how did you get to the subtraction?

OpenStudy (anonymous):

\[\ln{a \over b} = \ln a - \ln b \]

OpenStudy (anonymous):

oh of course!! i knew i was missing something

OpenStudy (anonymous):

(: